OpenStudy (samigupta8):

A piston can slide without friction inside a horizontal cylindrical vessel,which contains an ideal mono atomic gas.The piston and the cylinder both are made of a perfect heat insulating material. Initially the piston is in equilibrium and divides the cylinder into two parts A and B,which are not necessarily equal.The temperature of the gases in both the parts are equal.Now,the piston is held in its initial position and the gas in part A is supplied some amount of heat,then the piston is released.What will the piston do in its subsequent motion.

OpenStudy (samigupta8):

@ganeshie8 @Michele_laino @parthkohli

OpenStudy (samigupta8):

Sir I made some equations.. Shall I inform you about the same??

OpenStudy (michele_laino):

yes! thanks!

OpenStudy (samigupta8):

Firstly we know that moles will remain unchanged in part A and part B both so First is like Pa*r/Ta=PA*(r+x)/Ta' Here r is the distance of the piston from extreme (left) n I assumed the whole length of container as L X is the displacement of piston after supply of heat

OpenStudy (samigupta8):

Away from gas in part A...

OpenStudy (samigupta8):

2nd is Pb*(L-r)/Ta=Pb'(L-r-x)/Tb'

OpenStudy (samigupta8):

I took temperature in the first condition of vessel to be Ta

OpenStudy (michele_laino):

I ask, why do you think that there will be an expansion?

OpenStudy (samigupta8):

Heat is supplied to the part A that's why I thought so?? Am I wrong in my approach?

OpenStudy (michele_laino):

I made that question, since it is what the exercise is asking

OpenStudy (samigupta8):

Since it's an adiabatic case so v can say that the internal energy would be providing the necessary work for expansion

OpenStudy (michele_laino):

ok! yes, that's right! here is the equation for an adiabatic trasformation: \[\Large {P^{1 - k}}{T^k} = {\text{costant}}\]

OpenStudy (samigupta8):

Sir I got this now... Pa*r^y=Pa'*(r+x)^y And Pa*(L-r)^y=Pb'*(L-r-x)^y And y stands for the adiabatic constant(gamma)

OpenStudy (michele_laino):

where: \(k=C_p/C_v\) please wait, from the first principle of thermodynamic, we have this: \[\Large \Delta U = - L > 0 \Rightarrow L < 0\] in other words, the work done by th fluid A has to be negative

OpenStudy (michele_laino):

thermodynamics*

OpenStudy (samigupta8):

Sir i applied this concept PV^y =k

OpenStudy (michele_laino):

yes! It is the same

OpenStudy (samigupta8):

From that i got those two equation which i wrote just above ..

OpenStudy (michele_laino):

I saw your equations, please comment my last statement

OpenStudy (samigupta8):

And furthermore v can say that (Pa'-Pb')A=ma

OpenStudy (samigupta8):

I didn't see yours at the first attempt..! And yes sir they r definitely the same!!

OpenStudy (samigupta8):

M is mass of piston

OpenStudy (michele_laino):

please, from my last statement: \(L<0\) we can say that fluid A is doing a negative work, namely its volume will decrease, am I right?

OpenStudy (samigupta8):

Yep! It is ... Well i didn't give it a damn becoz what we r concerned here with is the motion of mass m ...so we can say that it will not change much to our ans

OpenStudy (samigupta8):

But if you wish the same then i can repicturise these equation..

OpenStudy (michele_laino):

I think that there is something wrong. Namely we can say that the fulid A is doing an expansion, then the first principle, \[\Delta U = - L > 0 \Rightarrow L < 0\] can be applied for fluid B only, so we can say that fluid B is doing a contraction and your equations are right!

OpenStudy (samigupta8):

Sir bt i m unable to make a shm type equation from those two particular equations... Will u pls help in that portion?

OpenStudy (michele_laino):



OpenStudy (michele_laino):

qualitatively spaking, we have solved the question using the first principle of thermodynamics: \[\large Q = \Delta {U_A} + {L_A},\quad {L_A} > 0\] for part A, and: \[\Large \Delta {U_B} = - {L_B} > 0 \Rightarrow {L_B} < 0\] for part B, since fluid inside B undergoes to an adiabatic compression

OpenStudy (michele_laino):

and the temperature of fluid B will increase

OpenStudy (samigupta8):

U mean to say that temperature in A will decrease! And that in B increase

OpenStudy (samigupta8):

Sir finally i got theses.. Pa[r^y(r+yx)-(L-r)^y(L-r+yx)]=ma

OpenStudy (michele_laino):

I mean that the temperature of B only, will increase

OpenStudy (samigupta8):

These*

OpenStudy (michele_laino):

do you want to write an equation of motion for piston ?

OpenStudy (samigupta8):

Sorry i made a mistake in very first part it should be r^y(r-yx)... Yes sir!!

OpenStudy (michele_laino):

please, do you know what is the type of motion of the piston?

OpenStudy (samigupta8):

I will know it after knowing the equation.

OpenStudy (michele_laino):

I think that it can be an oscillatory motion

OpenStudy (samigupta8):

If it comes out to be a is directly proportional to x then it is shm

OpenStudy (samigupta8):

How will v prove it?? I told the final equations that i came up with .. After that i m not getting ny way to solve them.. Will u like to help in these?

OpenStudy (michele_laino):

it is not an easy question, since there are many variables, namely temperatures of A and B, pressures of A and B

OpenStudy (samigupta8):

Sir bt the equation i made finally was free of all the variables and only there were Pa, r ,L and x which is great thing to solve the equation. .

OpenStudy (samigupta8):

The only thing is v have to bring in terms of x n then v r done!!

OpenStudy (michele_laino):

we start to give a little displacement to the piston, therefore we have to write the forces which are acting on such pistons. all what we know, is that after the heat exchage, we have all adiabatic trasformation of A and B fluids

OpenStudy (michele_laino):

piston*

OpenStudy (samigupta8):

I gave that x was the slight displacement only...

OpenStudy (michele_laino):

please wait, I'm writing some equations

OpenStudy (michele_laino):

after a little displacement \(x\) to the right, we get the new values of volume and temperature, as function of the old ones, like below: \[\Large \begin{gathered} {V_A}' = {V_A} + Sx \Rightarrow {P_A}' = {P_A}{\left( {\frac{{{V_A}}}{{{V_A} + Sx}}} \right)^k} \hfill \\ \hfill \\ {V_B}' = {V_B} - Sx \Rightarrow {P_B}' = {P_B}{\left( {\frac{{{V_B}}}{{{V_B} - Sx}}} \right)^k} \hfill \\ \end{gathered} \] where \(S\) is the cross sectional area of the piston.

OpenStudy (samigupta8):

This Va/Va+Sx can be even simplified to r/r+x...

OpenStudy (michele_laino):

so we can write this differential equation: \[\Large M\ddot x = \left\{ {{P_A}{{\left( {\frac{{{V_A}}}{{{V_A} + Sx}}} \right)}^k} - {P_B}{{\left( {\frac{{{V_B}}}{{{V_B} - Sx}}} \right)}^k}} \right\}S\] where \(M\) is the mass of the piston

OpenStudy (samigupta8):

r is the initial length of part A

OpenStudy (samigupta8):

Correct sir!!

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

now, as we can see, such ODE is not a \(linear\) ODE, so we have to linearize it

OpenStudy (samigupta8):

By binomial ...ryt?? I applied that even..after that only i got these.. Pa[r^y(r-yx)-(1-r^y)(L-r+yx)]=ma

OpenStudy (michele_laino):

\[\Large M\ddot x = \left\{ {{P_A}{{\left( {\frac{r}{{r + x}}} \right)}^k} - {P_B}{{\left( {\frac{{d - r}}{{d - r - x}}} \right)}^k}} \right\}S\]

OpenStudy (samigupta8):

Sure sir...these are in match vid my equation

OpenStudy (samigupta8):

Now if v bring that denominator term in numerator side v can solve them using binomial..

OpenStudy (michele_laino):

I think that we can write the Taylor's expansion at the first order, for both this functions: \[\Large {\left( {\frac{r}{{r + x}}} \right)^k},\quad {\left( {\frac{{d - r}}{{d - r - x}}} \right)^k}\]